5.Find the equation of line that passes through the points: A(2,3) and B(4,6)
a.2y=3x
b.−3x+2y=−2
c.3x+2y=−2
d.(3x+2y=6)

6.The equation of a straight line at point (-2,3) and slope 3 is given by
a.y=5x-3
b.y=3(x+3)
c.y=3(x-3)
d.y=3x-5
1
Expert's answer
2019-04-10T10:16:20-0400

5. The equation of line that passes through the points: A(2,3) and B(4,6) is: (y - 3)/3 = (x - 2)/2,

2y - 6 = 3x - 6,

2y = 3x, so the correct answer is a.

6.The equation of a straight line at point (-2,3) and slope 3 is given by the line, for which:

(y - y1)/(x - x1) = k,

(y - 3)/(x + 2) = 3,

y - 3 = 3x + 6,

y = 3x + 9 = 3(x + 3), so the correct answer is b.